OpenStudy (anonymous):

Evaluate the expression log(10,000x)

OpenStudy (anonymous):

here we go again...

OpenStudy (anonymous):

first step is to break it apart into a sum second step is to realize you know what \[\log(10,000)\] is because this means \[\log_{10}(10,000)\]

OpenStudy (anonymous):

let me know when you have the answer if this is not clear

OpenStudy (anonymous):

so solve logbase10 10,000?

OpenStudy (anonymous):

right, and log base ten is just a fancy way of saying "count the zeros" if you have a power of ten

OpenStudy (anonymous):

I got 4

OpenStudy (anonymous):

me too!

OpenStudy (anonymous):

ok now thats it?

OpenStudy (anonymous):

yeah that's it, although you started with \[\log(10,000x)=\log(10,000)+\log(x)=4+\log(x)\] like the last ones

OpenStudy (anonymous):

idk why its easier for someone to walk me through it